1.  y(4)+8y''+16y=0       ans, y=(c1+c2x)cos2x+(c3+c4x)sin2x
  2. y(4)+4y''+4y=0          ans, y=(c1+c2x)cos "sqr2" x+(c3+c4x)sin"sqr2" x
1
Expert's answer
2021-12-08T16:44:40-0500

1.

Corresponding (auxiliary) equation


"r^4+8r^2+16=0"

"(r^2+4)^2=0"

"r^2=-4"

"r=\pm 2i"

The general solution of the given differential equation is


"y=c_1\cos(2x)+c_2x\cos(2x)"

"+c_3\sin(2x)+c_4x\sin(2x)"

2.

Corresponding (auxiliary) equation


"r^4+4r^2+4=0"

"(r^2+2)^2=0"

"r^2=-2"

"r=\pm i\sqrt{2}"

The general solution of the given differential equation is



"y=c_1\cos(\sqrt{2}x)+c_2x\cos(\sqrt{2}x)"

"+c_3\sin(\sqrt{2}x)+c_4x\sin(\sqrt{2}x)"